OpenStudy (anonymous):

We decide to open a casino where we offer only one game. We win a dollar 60% of the time, and our customer wins a dollar 40% of the time. QUESTION: How much profit would we expect to have if 5,000 games are played?

OpenStudy (anonymous):

\[5000\times 0.6-5000\times 0.4=5000\times 0.2\]\]

OpenStudy (anonymous):

1,000 @satellite73

OpenStudy (anonymous):

yeah that looks right

OpenStudy (anonymous):

thank you i appreciate it

OpenStudy (anonymous):

yw